# DeFiScope: Detecting Various DeFi Price Manipulations with LLM Reasoning

### v1

@@ Abstract.
In this paper, we introduce the first LLM-based approach, DeFiScope, for detecting DeFi price manipulation attacks in both standard and custom price models. Our insight is that large language models (LLMs) have certain intelligence to abstract price calculation from code and infer the trend of token price changes based on the extracted price models. To further strengthen LLMs in this aspect, we leverage Foundry to synthesize on-chain data and use it to fine-tune a DeFi price-specific LLM. Together with the high-level DeFi operations recovered from low-level transaction data, DeFiScope detects various DeFi price manipulations according to systematically mined patterns. Experimental results show that DeFiScope achieves a high precision of 96% and a recall rate of 80%, significantly outperforming SOTA approaches. Moreover, we evaluate DeFiScope’s cost-effectiveness and demonstrate its practicality by helping our industry partner confirm 147 real-world price manipulation attacks, including discovering 81 previously unknown historical incidents.
copyright: none
1. Introduction
DeFi represents a form of finance that eliminates traditional financial intermediaries by utilizing smart contracts on a blockchain. A smart contract (Zou et al., 2019) is a self-executing program with the terms of the agreement between the buyer and the seller directly written into lines of code. The code and the agreements contained therein exist across a distributed, decentralized blockchain network (Dinh et al., 2018). Smart contracts enable trusted transactions and agreements to be carried out among disparate, anonymous parties without the need for a central authority, legal system, or external enforcement mechanism.
@@ 1. Introduction
While price manipulation is a pervasive risk in traditional financial markets, it poses even more significant damage in the context of DeFi. In the cryptocurrency space, attackers strategically add or remove large amounts of tokens in specific contracts, or impact the total supply of target tokens by burning or minting, to exploit vulnerabilities in the pricing mechanisms of DeFi applications, ultimately inflating or deflating the price of cryptocurrencies (Cointelegraph, 2024). To detect price manipulation attacks in DeFi applications, researchers have proposed several state-of-the-art (SOTA) approaches, primarily focusing on transaction monitoring-based methods that offer real-time protection, e.g., DeFiRanger (Wu et al., 2023) and DeFort (Xie et al., 2024).
@@
However, our analysis indicates that these SOTA approaches are less effective to non-standard price models in custom DeFi protocols. This is because they typically require token exchange rates to identify abnormal price changes, which are only suitable to calculate under standard price models such as CPMM (Constant Product Market Makers) and Stableswap Invariant (detailed in §2.2). Unfortunately, our subsequent evaluation shows that 44.2% of the 95 DeFi price manipulation attacks reported in the past three years used non-standard price models. Hence, instead of explicitly calculating the exchange rate for a pair of tokens, we aim to capture only the abnormal price fluctuations of tokens, which can be directly derived from the high-level price model and changes in token balances.
@@
To achieve this objective, we sought help from large language models (LLMs), considering that their trained intelligence might aid in inferring price changes associated with DeFi operations during the transaction process. As such, we introduce DeFiScope, the first LLM-based approach to detecting DeFi price manipulation attacks in both standard and custom price models. DeFiScope features several novel designs, such as constructing a transfer graph to recover high-level DeFi operations (§5) and systematically mining price manipulation patterns across all types of DeFi applications (§6). Among them, the key design is to fine-tune a DeFi price-specific LLM (§4), where we propose (i) simulating transactions using Foundry (Github, 2024) to generate our own fine-tuning data, and (ii) conducting a Chain-of-Thought (CoT)-style fine-tuning that integrates both on-chain data and the price context. The ablation study shows that fine-tuning increases the overall detection effectiveness by up to 31% compared to the baseline LLM (under DeFiScope) without fine-tuning.
@@
To thoroughly evaluate DeFiScope’s effectiveness and practicality, we conduct both a benchmark and a large-scale experiment. In the benchmark experiment, we collected 95 real-world price manipulation attacks from multiple sources and compared DeFiScope with three SOTA tools. The results show that DeFiScope achieves a superior detection accuracy (recall) of 80%, significantly higher than 51.6% with DeFiRanger, 52.6% with DeFort, and 35.8% with DeFiTainter (Kong et al., 2023) (a static source code analysis tool). In a large-scale experiment involving 968 suspicious transactions (a mix of various DeFi attacks, not just price manipulations) from our industry partner and 96,800 benign transactions from DeFort (Xie et al., 2024), we helped confirm 147 out of 153 price manipulation attacks reported by DeFiScope, yielding a precision of 96%. In particular, 81 of these attacks are previously unknown historical incidents. Moreover, DeFiScope achieves zero false alarms on benign transactions, with an average of 2.5-second time overhead across suspicious and benign transactions and merely \$0.0107 per-request LLM inference cost.
@@
Open Science and Ethics. To facilitate future research, all source code and experimental results are available in an (anonymous) GitHub repository (Anonymous, 2024). For ethical research, we ensure that all attacks elaborated in this work have been published or investigated by authoritative security companies through their established channels, and did not involve any undisclosed zero-day vulnerabilities.
2. Background and Motivation
As Figure 1 shows, on blockchains, two primary account types exist, i.e., externally owned accounts (EOAs), controlled by individuals using private keys, and contract accounts (CAs), governed by their contract code. External transactions are initiated externally by EOAs, while internal transactions occurs when a smart contract calls another contract internally. Smart contracts exchange information with other smart contracts through internal transactions, where a transaction involving communication between contracts can be regarded as a sequence of function calls, which has been greatly utilized in DeFi protocols for interoperability. For simplicity, we denote as user invocation an internal transaction between the user-controlled smart contracts and other smart contracts, which plays a key role in price manipulation attacks.
2.1. DeFi Protocols
A DeFi protocol typically consists of multiple smart contracts and offers various services including trading on decentralized exchange, lending, and yield farming.
Decentralized EXchange (DEX). A DEX is a peer-to-peer marketplace where users can trade cryptocurrencies. The two predominant forms of DEXs are order book DEXs and automated market makers (AMMs). In the order book DEXs, the management and matching processes of open buy and sell orders take place off-chain while the settlement of trades occurs on-chain. Alternatively, the AMM is fully on-chain, and maintains liquidity through liquidity pools without replying on real-time buyers and sellers. A liquidity pool maintains two or more tokens, and the prices of these tokens are determined by a constant mathematical formula. In this paper, we only target fully on-chain DEXs using AMM (Chainlink, 2024; coinbase, 2024).
Lending. To borrow assets from a lending protocol, borrowers are required to deposit other cryptocurrencies as collateral first. To ensure that borrowers can repay their loans, the collateral factor of supplied assets is utilized to calculate the maximum value of assets that users can borrow. For instance, consider a scenario where the collateral factor of ETH is set to 0.75, borrowers can only obtain assets up to 75% of the value of ETH they have deposited as collateral. If the value of the collateral decreases and falls below a certain level (the health factor of a borrower is less than 1), the liquidation of the collateral (the lending protocol sells the collateral at a discount rate) occurs to repay the borrowed amount (moai finance, 2024).
Yield Farming. It provides a way to earn rewards in the form of Annual Percentage Yields (APYs). A yield farmer could deposit cryptocurrencies into yield-farming applications, which invest these assets according to various strategies. Based on the returns from these investments, yield farmers can obtain rewards. There are two types of yield farming. The first involves users depositing cryptocurrencies into an application to obtain proof tokens, which can be traded, allowing holders to withdraw the corresponding deposited assets. The second, a stake-and-claim model, has users lock their assets into the contract and later claim them back. This process does not generate proof tokens, meaning the staked shares are non-tradable, and only the staking account can make the claim.
Figure 1. The blockchain transactions.
2.2. Price Models
@@
Constant Product Market Maker (CPMM). CPMM is one of the most prevalent DeFi AMMs and used in many well-known DEXs (Uniswap, 2024; PancakeSwap, 2024). It maintains liquidity through a constant function $R_{x}*R_{y}=k$, where $k$ is a constant, and $R_{x}$ and $R_{y}$ represent the reserves of two distinct assets $token_{x}$ and $token_{y}$ in a liquidity pool. In a swap operation, let $\Delta x$ amount of $token_{x}$ can exchange for $\Delta y$ amount of $token_{y}$, resulting in $(x+\Delta x)*(y-\Delta y)=k$ where the instantaneous price of $token_{x}$ denominated in $token_{y}$ is $P_{x,y}=\frac{y}{x}$ (Wu et al., 2023; Gogol et al., 2024).
Stableswap Invariant. The Stableswap Invariant (Curve.fi, 2024c), widely used in Curve AMMs (Curve.fi, 2024a), and is defined as follows:
@@
where $a$ indicates a constant amplification coefficient and n is the number of tokens in the liquidity pool, while $D$ represents the total underlying value, i.e., the total amount of tokens in the pool when token prices are equal, and $x_{i}$ denotes the current reserve of $token_{i}$. As $a$ approaches zero, the function behaves similarly to a CPMM.
Custom Price Model. Besides the aforementioned common price models, DeFi applications can customize their own pricing mechanisms, which are often more diverse and complicated. For example, UwULend (UwULend, 2024) offers lending service, permitting users to borrow sUSDe by depositing cryptocurrencies as collateral. To ensure that the user can repay the debt, it is necessary to examine whether the total value of sUSDe are below that of the collateral when lending it. In this process, UwULend calculate the price of sUSDe by the following custom price model, distinct from the aforementioned CPMM and Stableswap Invariant models:
(1) $\begin{split}P_{sUSDe}=median(\{IP_{USDe,Pool_{1}},\cdots,IP_{USDe,Pool_{5}},\\ EMAP_{USDe,Pool_{1}},\cdots,EMAP_{USDe,Pool_{5}}\})\end{split}$
@@
2.3. A Motivating Example
Figure 2. A simplified attack flow against UwULend.
Our approach is motivated by a real-world price manipulation attack on UwULend (Exvul, 2024; Etherscan, 2024a) in 2024. Figure 2 shows a attack flow of the price manipulation towards UwULend. The attack includes six key steps, and the root cause is the flawed price dependency in eq. 1 related to two functions borrow and liquidationCall.
@@ 2.3. A Motivating Example
By swapping a large amount of USDe into five liquidity pools, the attacker could deflate the instant price of USDe while the EMA prices remain unchanged, leading to a lower median price for sUSDe. Hence, the attack was able to borrow an exceedingly large amount of sUSDe. Subsequently, swapping back could increase the calculated price of sUSDe, resulting in the attacker’s collateral being unable to repay the debt , thereby allowing liquidation and acquiring deposited collateral with a bonus, leading to a \$19M loss.
@@
While detecting a price manipulation attack, existing tools generally define the exchange rate between two different tokens as their price. For example, the price of sUSDe could be calculated by dividing the amount of WETH deposited into the protocol at step 3 by the amount of sUSDe borrowed from it at step 4. Yet, defining abnormal price change range or even precisely capturing token prices is not a trivial task.
@@
  • DeFort (Xie et al., 2024) uses historical exchange rates, i.e., prices, between two tokens for computing a so-called normal fluctuation range. However, it may be error-prone as historical prices, especially for low-liquidity or newly created pools, could vary significantly, thus being insensitive to the detection of subtle price manipulations. For instance, in the UwULend attack, slight manipulations observed in sUSDe prices — decreasing by 4.2 % or increasing by 4.43 %, did not exceed the predefined bounds that evades the detection by DeFort illustrated in our evaluation in §7.
@@
  • DeFiRanger (Wu et al., 2023) detects abnormal price changes by tracing token exchange sequences within a transaction and comparing token exchange rates at different steps. However, completely tracking these sequences can be challenging for complex transactions. In the UwULend attack, the attacker crafted complicated deposit and withdrawal operations concerning WETH, making it difficult for DeFiRanger to detect.
@@
To address the aforementioned issues about price identification, our observation is that abnormal price fluctuations of tokens could be directly derived from the high-level price model and amount changes in token balances, without needing to explicitly calculate the exchange rate for a pair of tokens. Taking the UwULend attack as an example, swapping USDe into liquidity pools in step 2 decreases the value of $IP_{USDe,Pool_{i}}$, causing an abnormal drop in the median price, namely the price of sUSDe. Based on this observation, in this paper, we propose a novel price change reasoning approach powered by LLM and integrate with predefined rules to enhance the capability of price manipulation attack detection.
Figure 3. A high-level overview of DeFiScope.
3. Overview
@@
Based on the analysis of the motivating example illustrated in §2.3, we establish an intuition of inferring price change associated with DeFi operations during the transaction process to detect DeFi price manipulations in various scenarios. While inferring through standard price models (e.g., CPMM and Stableswap as introduced in §2.2) is straightforward, it is challenging to (i) interpret low-level price calculation Solidity code into high-level price calculation formulas, and (ii) infer the price change from the (custom) price calculation formulas and the information on the token balance changes in related accounts. Both tasks require certain intelligence.
@@
To address this key challenge, we introduce the first LLM-based approach, DeFiScope, for effective DeFi price manipulation detection. As depicted in Figure 3, DeFiScope consists of ten steps. In step ①, DeFiScope first decodes and slices raw transaction data. Then in steps ② ⑤ DeFiScope extracts the code of possible price calculation functions from smart contracts based on their signature, and token balance changes in relevant accounts. Subsequently, in steps ⑥, DeFiScope embeds these two information into a prompt template, which will be used by the LLM to extract the price model and infer the price change in steps ⑦ ⑧. In the meantime, in steps ③ ④, DeFiScope constructs the transfer graph (§5.1) and uses it to recover the high-level DeFi operations (§5.2) associated with those price change. This is because using the trend of token price changes alone is insufficient for detecting price manipulation. Finally, based on the recovered DeFi operations and their price change information, DeFiScope maps them into eight attack patterns listed in §6 and detects DeFi manipulation attacks in step ⑨.
@@
We conduct an off-line fine-tuning step (step ⓪) to enhance LLMs’ capabilities in extracting price calculation models and token price change reasoning. This is because while the off-the-shelf LLMs may exhibit certain capability in reasoning and code understanding, they are limited for predicting the trend of token price changes given unlabeled code snippet and numerical changes of tokens balance, which will be illustrated in §7. We will detail our fine-tuning technique in §4.
4. Price Change Inference with LLMs
Instead of leveraging traditional symbolic or concrete execution techniques to estimate the price model through input and output data points, which results in a significant gap between the actual price calculation formulas and the approximating ones, we use LLMs to understand codes calculating token prices and extract corresponding price models, which eliminates the gap caused by estimation.
4.1. LLM Fine-tuning

### v2

# Detecting Various DeFi Price Manipulations with LLM Reasoning

@@ Abstract
In this paper, we introduce the first LLM-based approach, DeFiScope, for detecting DeFi price manipulation attacks in both standard and custom price models. Our insight is that large language models (LLMs) have certain intelligence to abstract price calculation from smart contract source code and infer the trend of token price changes based on the extracted price models. To further strengthen LLMs in this aspect, we leverage Foundry to synthesize on-chain data and use it to fine-tune a DeFi price-specific LLM. Together with the high-level DeFi operations recovered from low-level transaction data, DeFiScope detects various DeFi price manipulations according to systematically mined patterns. Experimental results show that DeFiScope achieves a high recall of 80% on real-world attacks, a precision of 96% on suspicious transactions, and zero false alarms on benign transactions, significantly outperforming SOTA approaches. Moreover, we evaluate DeFiScope’s cost-effectiveness and demonstrate its practicality by helping our industry partner confirm 147 real-world price manipulation attacks, including discovering 81 previously unknown historical incidents.
I Introduction
@@
DeFi represents a form of finance that eliminates traditional intermediaries by utilizing smart contracts on a blockchain. A smart contract [1] is a self-executing program with the terms of the agreement directly written into code, deployed on a distributed, decentralized blockchain network [2]. These contracts power decentralized applications such as Automated Market Makers (AMMs), where token prices are determined by on-chain pricing mechanisms rather than centralized order books. While price manipulation is a pervasive risk in traditional finance, DeFi pricing mechanisms create more attack surfaces: adversaries can strategically add or remove liquidity, or impact token supply by minting or burning, to exploit vulnerabilities in pricing mechanisms, ultimately distorting prices of cryptocurrencies [3]. To detect price manipulation attacks in DeFi applications, researchers have proposed several state-of-the-art (SOTA) approaches, primarily focusing on transaction monitoring-based methods that offer real-time protection, e.g., DeFiRanger [4] and DeFort [5].
@@
However, our analysis indicates that these SOTA approaches are less effective to non-standard price models in custom DeFi protocols. This is because they typically require token exchange rates to identify abnormal price changes, which are only suitable to calculate under standard price models such as CPMM (Constant Product Market Makers) and Stableswap Invariant (detailed in §II-A). Unfortunately, our subsequent evaluation shows that 44.2% of the 95 DeFi price manipulation attacks reported in the past three years used non-standard price models. Hence, instead of explicitly calculating the exchange rate for a pair of tokens, we aim to capture only the abnormal price fluctuations of tokens, which can be directly derived from the high-level price model and changes in token balances.
@@
To achieve this objective, we sought help from large language models (LLMs), considering that their trained intelligence might aid in inferring price changes associated with DeFi operations during the transaction process. As such, we introduce DeFiScope, a source code-level detector and the first LLM-based approach to detecting DeFi price manipulation attacks in both standard and custom price models. DeFiScope features several novel designs, such as constructing a transfer graph to recover high-level DeFi operations (§V) and systematically mining price manipulation patterns across four prevalent types of DeFi applications (§VI). Among them, the key design is to fine-tune a DeFi price-specific LLM (§IV), where we propose (i) simulating transactions using Foundry [6] to generate our own fine-tuning data, and (ii) conducting a Chain-of-Thought (CoT)-style fine-tuning that integrates both on-chain data and the price context. The ablation study shows that fine-tuning increases the overall detection effectiveness by up to 31% compared to the baseline LLM (under DeFiScope) without fine-tuning.
@@
To thoroughly evaluate DeFiScope’s effectiveness and practicality, we conduct both a benchmark and a large-scale experiment. In the benchmark experiment, we collected 95 real-world price manipulation attacks from multiple sources and compared DeFiScope with three SOTA tools. The results show that DeFiScope achieves a superior detection accuracy (recall) of 80%, significantly higher than 51.6% with DeFiRanger, 52.6% with DeFort, and 35.8% with DeFiTainter [7] (a static source code analysis tool). In our large-scale experiment, we evaluated DeFiScope on two datasets provided by our industry partner and DeFort [5]. The first dataset consists of 968 suspicious transactions, covering a variety of DeFi attacks beyond price manipulations. On this dataset, DeFiScope flagged 153 potential price manipulation attacks, 147 of which were manually verified as true positives, yielding a precision of 96%. Notably, 81 of these attacks are previously undocumented price manipulation incidents. The second dataset involves 96,800 benign transactions, on which DeFiScope produced zero false alarms. Regarding efficiency, DeFiScope incurs an average time overhead of 2.5 seconds per transaction across both suspicious and benign transactions, with a per-request LLM inference cost of only \$0.0107.
@@
Artifact. To facilitate open science and future research, all source code, experimental results, and supplementary material are available in https://github.com/AIS2Lab/DeFiScope.
II Background and Motivation
On blockchains, two primary account types exist, i.e., externally owned accounts (EOAs), controlled by individuals using private keys, and contract accounts (CAs), governed by their contract code. External transactions are initiated externally by EOAs, while internal transactions occurs when a smart contract calls another contract internally. Smart contracts exchange information with other smart contracts through internal transactions, where a transaction involving communication between contracts can be regarded as a sequence of function calls, utilized by DeFi protocols for interoperability. For simplicity, we denote as user invocation an internal transaction between the user-controlled smart contracts and other smart contracts, which plays a key role in price manipulation attacks.
II-A Price Models
@@
Constant Product Market Maker (CPMM). CPMM is one of the most prevalent DeFi AMMs and used in many well-known DEXs [8, 9]. It maintains liquidity through a constant function xy=kx*y=k, where $k$ is a constant, and $x$ and $y$ represent the reserves of two distinct assets $token_{x}$ and $token_{y}$ in a liquidity pool. In a swap operation, let $\Delta x$ amount of $token_{x}$ can exchange for $\Delta y$ amount of $token_{y}$, resulting in $(x+\Delta x)*(y-\Delta y)=k$ where the instantaneous price of $token_{x}$ denominated in $token_{y}$ is $P_{x,y}=\frac{y}{x}$ [4, 10].
Stableswap Invariant. The Stableswap Invariant [11], widely used in Curve AMMs [12], and is defined as follows:
@@
where $a$ indicates a constant amplification coefficient and $n$ is the number of tokens in the liquidity pool, while $D$ represents the total amount of tokens in the pool when token prices are equal, and $x_{i}$ denotes the current reserve of $token_{i}$. The curve is nearly flat in the middle part, while price changes at the extremes resemble CPMM behavior.
Custom Price Model. Besides the mentioned common price models, DeFi protocols can customize their own pricing mechanisms, which are often more diverse and complicated. For example, UwULend [13] offers lending service, permitting users to borrow sUSDe by depositing cryptocurrencies as collateral. To ensure the user can repay, examining whether the total value of sUSDe are below that of the collateral when lending it is necessary. In this process, UwULend calculates the price of sUSDe by the following custom price model, distinct from the aforementioned CPMM and Stableswap Invariant models:
$\begin{split}P_{sUSDe}=median(\{IP_{USDe,Pool_{1}},\cdots,IP_{USDe,Pool_{5}},\\ EMAP_{USDe,Pool_{1}},\cdots,EMAP_{USDe,Pool_{5}}\})\end{split}$ (1)
@@
II-B A Motivating Example
@@
Our approach is motivated by a real-world price manipulation attack on UwULend [20, 21] in 2024. The root cause of the attack is the flawed price dependency in eq. 1 related to two functions borrow and liquidationCall. By swapping a large amount of USDe into five liquidity pools, the attacker deflated the instant price of USDe while the EMA prices kept unchanged, resulting in a lower median price for sUSDe. Hence, the attack was able to borrow an exceedingly large amount of sUSDe using the same amount of collateral. Then, swapping back could increase the calculated price of sUSDe, resulting in the attacker’s collateral being unable to repay the debt, thereby allowing liquidation and acquiring deposited collateral with a bonus, leading to a \$19M loss.
@@
While detecting a price manipulation attack, existing tools generally define the exchange rate between two different tokens as their price. For example, the price of sUSDe could be calculated by dividing the amount of WETH deposited into the protocol by the amount of sUSDe borrowed from it. Yet, defining abnormal price change range or even precisely capturing token prices is not a trivial task.
Figure 1: A high-level overview of DeFiScope.
@@
  • DeFort [5] uses historical exchange rates, i.e., prices, between two tokens to compute a so-called normal fluctuation range. However, it may be error-prone as historical prices, especially for low-liquidity or newly created pools, could vary significantly, thus being insensitive to the detection of subtle price manipulations. For instance, in the UwULend attack, slight manipulations observed in sUSDe prices — decreasing by 4.2 % or increasing by 4.43 %, did not exceed the predefined bounds that evades DeFort’s detection, illustrated in our evaluation in §VII.
@@
  • DeFiRanger [4] detects abnormal price changes by tracing token exchange sequences within a transaction and comparing token exchange rates. However, completely tracing these sequences is challenging for complex transactions. In the UwULend attack, the attacker crafted complicated deposit and withdrawal operations concerning WETH, making it difficult for DeFiRanger to detect.
@@
To address the aforementioned issues about price identification, our observation is that abnormal price fluctuations of tokens could be directly derived from the high-level price model and amount changes in token balances, without needing to explicitly calculate the exchange rate for a pair of tokens. Taking the UwULend attack as an example, swapping USDe into liquidity pools decreases the value of $IP_{USDe,Pool_{i}}$, causing an abnormal drop in the median price, namely the price of sUSDe. Based on this observation, in this paper, we propose a novel price change reasoning approach powered by LLM and integrate with predefined rules to enhance the capability of price manipulation attack detection.
III Overview
@@
Based on the analysis of the motivating example illustrated in §II-B, we establish an intuition of inferring price change associated with DeFi operations during the transaction process to detect DeFi price manipulations in various scenarios. While inferring through standard price models (i.e., CPMM and Stableswap as introduced in §II-A) is straightforward, it is challenging to (i) interpret low-level price calculation Solidity code into high-level price calculation formulas, and (ii) infer the price change from the (custom) price calculation formulas and the information on the token balance changes in related accounts. Both tasks require certain intelligence.
@@
To address this key challenge, we introduce the first LLM-based approach, DeFiScope, for effective DeFi price manipulation detection. As depicted in Fig. 1, DeFiScope consists of ten steps. In step ①, DeFiScope first decodes and slices raw transaction data, and retrieves the source code of smart contracts called within the transaction. Then in steps ② ⑤ DeFiScope extracts the code of possible price calculation functions from smart contracts based on their signature, and token balance changes in relevant accounts. Subsequently, in steps ⑥, DeFiScope embeds these two information into a prompt template, which will be used by the LLM to extract the price model and infer the price change in steps ⑦ ⑧. In the meantime, in steps ③ ④, DeFiScope constructs the transfer graph (§V-A) and uses it to recover the high-level DeFi operations (§V-B) associated with those price change. This is because using the trend of token price changes alone is insufficient for detecting price manipulation. Finally, based on the recovered DeFi operations and their price change information, DeFiScope maps them into eight attack patterns listed in §VI and detects DeFi manipulation attacks in step ⑨.
@@
We conduct an off-line fine-tuning step (step ⓪) to enhance LLMs’ capabilities in extracting price calculation models and token price change reasoning. This is because while the off-the-shelf LLMs may exhibit certain capability in reasoning and code understanding, they are limited for predicting the trend of token price changes given unlabeled code snippet and numerical changes of tokens balance, which will be illustrated in §VII-B. We now detail our fine-tuning technique in §IV.
IV Price Change Inference with LLMs
IV-A LLM Fine-tuning